Calculate the pH and the pOH of an aqueous solution that is 0.050 M in HCl(aq) and 0.085 M in HBr(aq) at 25 °C. pH = pOH =
Answer #1

PAGE: DATE: 11 Answer We want to calcuate the pH and the post of an aqueous solution that is 0.050 Min Hel caq) and 0.085 M HPAGE: DATE: 11 PH= 0.8696 we know that, pH + POH = 14 !! ! pOH = 14-PH pOH = 14-0.8696 pOH = 13.1303:... ., pOH = 13.1303 PH